Wegner estimate with local continuity requirements on the coupling constants

Ivan Veselic Ruhr-Universität Bochum, Germany

(joint work with W. Kirsch)

The talk concerns spectral properties of random Scheödinger operators. We prove a Wegner estimate for alloy type models near the infimum of the spectrum. The single site potential is assumed to be non-positive and the coupling constants non-negative. The distribution of a single coupling constant needs to posses a bounded density only near the upper edge of its support.